The compound KClO3 decomposes according to the following equation:


a. What is the mole ratio of KClO3 to O2 in this reaction: The mole ratio KCLO3 to O2 is 2 : 3

b. How many moles of O2 can be produced by letting 6.0 moles of KClO3 react based on the above equation? n (O2) = 3/2 x n (KClO3) = 3/2 x 6 = 9 moles

c. How many molecules of oxygen gas, O2, are produced in question  1b? 1 mole = 6.022 x 1023 molecules 

9 moles x 6.022 x 1023 molecules = 54.198 ·1023molecules

2. Magnesium combines with chlorine, Cl2, to form magnesium chloride, MgCl2, during a synthesis reaction.

a. Write a balanced chemical equation for the reaction. Mg + Cl2= MgCl2

n (MgCl2) = n (Cl2)


b. How many moles of magnesium chloride can be produced with 3 moles of chlorine? 

3 moles of Cl2 gives 3 moles of MgCl2

3. Ethanol burns according to the following equation. If 15.0 mol of C2H5OH burns this way, how many moles of oxygen are needed?


4. Solutions of iron (III) chloride, FeCl3, are used in photoengraving and to make ink. This compound can be made by the following reaction:


a.  How many grams of Fe are needed to combine with 4.5 moles of Cl2?

b. If 240 g of Fe is to be used in this reaction, with adequate Cl2, how many moles of FeCl3 will be produced?


5. Ammonia is produced synthetically by the reaction below. How many moles of NH3 are formed when 200.0 g of N2 reacts with hydrogen?


6. Iron metal is produced in a blast furnace by the reaction of iron (III) oxide and coke (pure carbon). If 25.0 moles of pure Fe2O3 is used, how many grams of iron can be produced? The balanced chemical equation for the reaction is:


7. Aluminum oxide is decomposed using electricity to produce aluminum metal. How many grams of aluminum metal can be produced from 100.0 g of Al2O3?


8. Predict the mass of hydrogen gas produced when 2.73 g of aluminum reacts in a single replacement reaction with excess sulphuric acid.





1
Expert's answer
2021-01-19T04:29:58-0500

a) The mole ratio KCLO3 to O2 is 2 : 3.

b) n (O2) = 3/2 x n (KClO3) = 3/2 x 6 = 9 moles.

c) 1 mole = 6.022 x 1023 molecules 9 moles x 6.022 x 1023 molecules = 54.198 ·1023molecules

2. a) Mg + Cl2= MgCl2

n (MgCl2) = n (Cl2).

b) 3 moles of Cl2 gives 3 moles of MgCl2.

3) 45 mol.

4) a) 2Fe + 3Cl2 =2 FeCl3

2-3

X-4.5 x= 3.

b) 240 x 162.5/56= 696.42.

5) 34 x 200/28 = 242.85.

6) 25 x 2 = 50 moles.

7) 100 x 54 /102= 52.94 g.

8) 2.73 x 6 / 54= 0.3 g.